Find the equation of the line which passes through the point (-3,5) and is perpendicular to the line 4x+3y=6. Express your answer in slope-intercept form.
Found 4 solutions by Alan3354, stanbon, ikleyn, amalm06:
Answer by Alan3354(69443)

Been done 1000's of times. Look it up.

Answer by stanbon(75887)

Find the equation of the line which passes through the point (-3,5) and is perpendicular to the line 4x+3y=6. Express your answer in slope-intercept form.
Find the slope of the given equation.
y = (-4/3)x + 2
slope = -4/3
-----
Find the slope of a line that is perpendicualar to the given line.
m = (3/4)
------
Ans Form:: y = mx + b
Solve for "b" if m = 3/4, x = -3, y = 5.
5 = (3/4)(-3) + b
5 = -9/4 + b
29/4 = b
-----
Ans:: y = (3/4)x + (29/4)
--------------
Cheers,
Stan H.
-------------

Answer by ikleyn(52783)

.
If you want to learn the subject ONCE FOR ALL, then read the lesson

    - Equation for a straight line perpendicular to a given line and passing through a given point

in this site and find there the complete explanation on how to do it.



If, in opposite, your task is simply  "copy,  paste,  submit and collect the answers",  then you may ignore this post.



Answer by amalm06(224)

You can use calculus to find the slope.
4x+3y=6

3y=6-4x

y=(-4/3)x+2

Differentiate implicitly:

dy=(-4/3)dx

dy/dx=-4/3

The slope of the line in question must be 3/4, since the product of the slopes of two perpendicular lines is -1.

To find the y-intercept of the line in question, note that the line passes through (-3,5). So that
5=(3/4)(-3)+b

b=29/4

Therefore, the equation of the line in question is

y=(3/4)x+(29/4)